x-2y=-1 and -x+5y=4
Found 2 solutions by Imasmart, richwmiller:
Answer by Imasmart(2)

line up the equations
x-2y=-1
-x+5y= 4
the x variable cancels out and your left with
-2y=-1
5y=4
then you subtract the to y variables and the to whole numbers and get
3y=3
then you divide by 3 on both sides and get
y=1
then you plug in 1 for the top equation
x-2(1)=-1
x-2=-1
add 2 on both sides and get the value of x
x=1

Your answer is x=1 and y=1
PS you can plug in the two values into either equation and check your answer

Answer by richwmiller(9143)

x-2y=-1
-x+5y=4
add
3y=3
y=1
-x+5=4
x=1
checks
1-2=-1
-1=-1
ok
-1+5=4
4=4